# Bending-strain effects in conventional superconductors and superconducting junctions

### v1

@@ II.3 Discretization and diagonalization
$-\frac{i}{2}\sum_{\langle\bm{i},\bm{j}\rangle,\sigma,\sigma^{\prime}}\alpha_{N}\>\hat{\mathcal{N}}(s)\cdot\left(\bm{\sigma}(s)\times\bm{d_{ij}}\right),$
@@
II.4 Order parameters and observables
To compare different symmetries, we define the superconducting order parameters as $\Delta_{\bm{i}}(\tau)=\langle\mathcal{T}c_{\bm{i},\sigma}(\tau)\>c_{\bm{i}^{\prime},\sigma^{\prime}}(0)\rangle$, where $\mathcal{T}$ is the time ordering operator, and $\tau$ is the relative time coordinate scaled by $\hbar/t$. Thus, we can investigate the order parameters of even- ($\tau=0$) and odd-frequency ($\tau\neq 0$) correlations. In Table 1, we introduce the notation for the relevant combinations of underlying Spin-Parity-Orbital-Time (SPOT) symmetries that ensure overall odd order parameter symmetry Linder and Balatsky (2019).
$S$ $P^{*}$ $O$ $T^{*}$
$\mathcal{S}_{0}$1 1 1 1 even-frequency $s$-wave singlet
$\mathcal{S}_{\sigma\sigma^{\prime}}(\tau)$ 1 1 1 -1 odd-frequency $s$-wave triplets
$\mathcal{P}^{\bm{m}}_{\sigma\sigma^{\prime}}$ 1 -1 1 1 even-frequency $p^{m}$-wave triplets
$\mathcal{P}^{\bm{m}}_{0}(\tau)$1 -1 1 -1 odd-frequency $p^{m}$-wave singlet
Here $\bm{m}$ is one lattice spacing in either the tangential or binormal direction, and the time-dependent operators are given by $c_{\bm{i},\sigma}(\tau)=e^{i\mathcal{H}\tau}c_{\bm{i},\sigma}e^{-i\mathcal{H}\tau}$. Observe here that the p-wave correlations have $\bm{i}-\bm{i}^{\prime}\neq 0$, and the s-wave correlations have $\bm{i}-\bm{i}^{\prime}=0$.
The gap equation in the mean-field approximation is given by $\Delta_{i}=U_{i}\>\mathcal{S}_{i,0}$, which in the new quasiparticle basis yields
where we have used that ${\big{<}{\gamma^{\dagger}_{n,k}\gamma_{n^{\prime},k}}\big{>}=f(2\mathcal{E}_{n,k})\delta_{n,n^{\prime}}}$ and $f(2\mathcal{E}_{n,k})$ is the Fermi-Dirac distribution, and $\beta$ is the inverse temperature. The corresponding expressions for Eq. 22-(26) in the quasiparticle basis of Eq. 21 that are used for computation are given in Appendix A for convenience.
$\sum_{m}\bm{J}_{\bm{i},m}=-i\big{[}\mathcal{H},\>\bm{S}_{\bm{i}}\big{]}.$ (29)
Here, $\mathcal{H}$ is the Hamiltonian and $m$ enumerates the surface normals of the unit cell $\bm{i}$. By comparing the current flowing in and out from a single unit cell in the tangential direction, a final expression can be obtained Risinggård and Linder (2019). Thus, $\bm{J}$ denotes the spin-current flowing parallel to $\hat{\mathcal{T}}$ in real space, with components in spin space $J^{\mu}$, where $\mu=(s,n,b)$ indicates spin alignment in the tangential, normal, and binormal directions. The components of the spin-current $\bm{J}$ are only conserved in areas without spin-orbit coupling. We will therefore only examine spin-current as a useful observable in cases where the material is not curved at $i$, and $\alpha_{N}\neq 0$, i.e. in an uncurved normal metal sandwiched between two curved superconductors, as in Fig. 1. By using the quasiparticle operators and elements defined in Eq. 17, we find the tangential- and binormal-components of the spin-current
III Results
We examine the effects of bending-strain in clean, conventional superconducting films in Section III.1. In Section III.2, we consider the additional effect of proximity-coupling the strained superconductor to a normal metal, and in Section III.3 we consider the proximity-induced effects in SNS junctions.
In the following, we normalize the curvature to the superconducting coherence length $\alpha_{N}=\kappa\xi$, with proportionality constant $a_{N}=1$. This can be used to relate the effective spin-orbit strength to the strain via $\alpha_{N}=\epsilon_{ss}\xi/n$, where $n=(N_{n}-1)a/2$ is the length from the center of the film to the outermost layer (see Fig. 1), and $a$ is the lattice spacing. Using this, any specific strain values indicated from now on will refer to the strain at the upper/lower plane of the superconductor. For illustrative purposes, we select parameters corresponding to superconducting niobium, with $\alpha_{N}=2.3\epsilon_{ss}/(N_{n}-1)$, where $\epsilon_{ss}$ is the strain in percent. The choice of material is arbitrary, chosen here to provide a familiar scale in comparison with purely spin-orbit-coupled systems, and the results presented here will apply to any conventional thin-film superconductorBlatt and Thompson (1963); Guo et al. (2004); Chiang (2004); Ummarino and Zaccone (2024). We examine film thicknesses $N_{n}\in(3,6,9)$, and set $\xi/a=2.3$ so that $\alpha_{N}$ falls within the common range for SOC-systems. We denote the total number of lattice sites in the tangential, normal, and binormal direction with $N_{s},N_{b}$, and $N_{b}$, respectively. The number of lattice sites in the tangential direction in the superconductors and normal metal individually will be denoted $N_{S}$ and $N_{N}$ respectively, as indicated in Fig. 2. The thin-film Hamiltonian is solved by summing over one-dimensional chains with a binormal momentum index $k$ and averaging over the first Brillouin zone, as described in Eqs. 14 and 20. Strain and curvature are incorporated as spin-orbit coupling, with the Pauli matrices rotating according to Eq. 4. The coordinate system is oriented such that the normal metal, when present, lies in the $y-z$ plane, as illustrated in Fig. 1.
@@ III.1 Bending-strain in a thin-film superconductor
We initialize a superconductor with an s-wave gap on a square lattice, which is solved self-consistently with an iterative diagonalization by updating the gap equation (15) with the expression in Eq. 27 until their absolute difference falls below a threshold, here set to $10^{-6}$. We set the number of sites in the binormal and tangential direction to be square, with $N_{b}=N_{s}=99$ in this subsection. By selecting an odd number of lattice sites, the end of the Brillouin zone does not contribute, i.e. we can ignore the last term in Eq. 20, since $ka=2\pi m/N_{b}$, where $m$ is an integer. While a specific number is chosen for illustration, the results are not sensitive to the number of lattice sites and remain consistent across different configurations, although a more sophisticated model would be expected to show atomic-scale oscillation of the strength of the superconductivity Guo et al. (2004); Ummarino and Zaccone (2024). Furthermore, normalizing to the hopping $t$, we set the attractive potential parameter to $U=1.5$, the chemical potential to $\mu=1$, and the temperature to $T=0.2$.
The direction of asymmetry of the strain-induced spin-orbit coupling in the superconductor is in $\hat{\mathcal{N}}(s)$, which creates a mixed parity s- and p-wave order parameter in the superconductor. The magnitude of the s-wave gap will diminish as a function of the strain, as shown in Fig. 3. This is analogous to doing self-consistency calculations for an increasing Rashba coefficient in a spin-orbit coupled superconductor, which yields the same results Ying et al. (2017).
We denote the p-wave triplet pairings with no spin projection as $\mathcal{P}^{s}_{\uparrow\downarrow}$, and the spin-polarized triplets as $\mathcal{P}^{b}_{\sigma}$, as indicated in Table 1. With time-inversion symmetry conserved, the spin-polarized p-wave amplitudes are equal in magnitude $|\mathcal{P}^{b}_{\uparrow}|=|\mathcal{P}^{b}_{\downarrow}|$. The superconducting gap closes completely for high enough strain, so that both s- and p-wave correlations become negligible. The thicker the film, the higher strain it can support while maintaining superconductivity. The curvature at a given strain is less for thicker films, and so the effective SOC is weaker. The singlet s-wave and finite triplet p-wave order pairings are shown in Fig. 3 as a function of the strain. The pairings arise through the entire superconductor because the superconductor has uniform circular curvature, giving a constant bending-strain profile and equivalent SOC-strength. If the strain is only applied to a certain part, the p-wave amplitudes will only be finite locally. For instance, in an ellipse, the correlations are localized at the apexesHeinrich (2024).
III.2 SN bilayer with bending-strain in S
Next, we consider a superconductor-normal metal (SN) bilayer, with an example of arbitrarily chosen length of superconductor with $N_{S}=50$ sites, a normal metal with $N_{N}=21$ sites, and $N_{b}=71$ sites in the binormal direction. The odd number here again simplifies the sum in Eq. 20, as the end of the Brillouin zone does not contribute, as in Section III.1. The film thickness in the normal direction is in this case set to $N_{n}=3$, and the strain is only applied to the superconductor.
In Fig. 4, we show the even-frequency pairing amplitudes and the relative time derivatives of the finite odd-frequency pairings at the interface. The even-frequency amplitudes originate from the bending-strain via the spin-asymmetry introduced by the effective spin-orbit coupling in the superconductor, while the interface between the materials induces the odd-frequency pairings. The even-frequency p-wave triplets $\mathcal{P}^{s}_{\uparrow\downarrow}$ persist throughout the bulk, and these triplets have an interface-dependence on both Andreev and normal reflections Cayao and Black-Schaffer (2018). They vanish for $\alpha_{N}=0$, in contrast to $\mathcal{P}^{s}_{0}(\tau)$, which remains finite, as can be seen in Fig. 4. Furthermore, since the superconductors are strained and the normal metal is not, the bending-induced spin-orbit coupling vanishes at the interfaces. This interfacial change acts on spins like a local spin splitting in the binormal direction, giving rise to an odd-frequency $\mathcal{S}_{\uparrow\downarrow}(\tau)$-wave triplet and $\mathcal{P}^{b}_{0}(\tau)$-wave singlet contribution.
Figure 4: The pairing amplitudes around the SN interface for increasing strain in the superconductor. The first column shows the non-zero even-frequency pairings and the second column is the relative time derivative of the odd-frequency pairings at $\tau=0$. The spin-polarized p-wave pairings are equal in magnitude $|\mathcal{P}^{b}_{\sigma}|=|\mathcal{P}^{b}_{-\sigma}|$. The derivatives are compared to the amplitudes of the corresponding odd-frequency pairings in Fig. B.1. System parameters are given in the main text.
For completeness, we compare the derivatives in Fig. 4 with the magnitudes of the corresponding odd-frequency pairings in Fig. B.1 in the Appendix. There we see that the amplitudes – which remain within the same order for relative times within that order – are comparable to the derivative evaluated at relative time $\tau=0$.
III.3 SNS junction with bending-strain in S
In Fig. 5, we show the tangential and binormal components of the spin current $\bm{J}$ at the center of the normal metal. The sign of $J^{b}$ depends on the length of the normal metal and strain, resulting in so-called $0-\pi$ oscillations as a function of both length and the effective spin-orbit coupling strength experienced, shown here as a function of the bending-strain in the superconductor. The $0-\pi$ oscillations in $J^{b}$ as a function of length are included in Appendix C for completeness. Changes in current-direction as indicated by $0-\pi$ oscillations are well known in Josephson junctions with a ferromagnetic weak link Ryazanov et al. (2001); Kontos et al. (2002); Birge and Satchell (2024), but in SNS junctions they are so-far only found when there exists a non-equilibrium electron distribution in N Baselmans et al. (1999). This result therefore indicates that, with regards to generating a Josephson current, strain, or effective SOC, acts on the normal metal in a similar way to a Zeeman interaction in a ferromagnet. A subgap spin Josephson effect has also been suggested in the topologically nontrivial phase of topological Josephson junctions with a magnetic field and SOC in the superconductors Mao and Sun (2022), and a superconducting diode effect has been noted due to a combination of SOC and magnetic field or intrinsically p-wave superconductors Mao et al. (2024). In our case, we show that this current can appear when using conventional s-wave superconductors with bending-strain and a simple normal metal weak link.
In Fig. 6, we show the spin-current $J^{s}$ again, this time as a function of both strain and position in $N$, alongside the induced spin-magnetization $S^{b}$. A combination of SOC and charge current in normal metals generically leads to spin accumulation Manchon et al. (2015); the charge current shifts the Fermi contour along the $k_{s}$-axis in momentum space, which, in the presence of SOC, creates a finite spin magnetization in the binormal direction. The appearance of a spin magnetization can therefore be attributed to the Edelstein effect Edelstein (1990). Since we have fixed the phases in the superconductors, the spin accumulation is strongest within the normal metal, as seen in Fig. 6. In our numerical scheme, we split the phase difference equally between the two superconductors. The diagonalization gives proximity-induced pairings in the normal metal, with a phase gradient $\partial_{s}\phi$ in the tangential direction. When this phase gradient is nonzero, a charge current will flow. This gradient diminishes away from the interfaces within the superconductors, so the spin magnetization [$S^{b}$] and charge current will eventually decay there.
We see from Fig. 6 that the switch in magnetization can be determined from the $0-\pi$ switching profile of $J^{s}$. The principal reason why no magnetic field or intrinsic p-wave order is required in this case is due to the combination of effective-SOC and interfacial effects on rotating the proximity-induced correlations, and is inherently linked to the thin-film nature of the system, which retains an additional term in the cross-product of Section II.3 compared with the one-dimensional analysis of most triplet transport analysis to date. When considering two dimensions, we get a $\sigma_{T}k_{s}$-term in the $sb$-plane in the presence of a phase gradient. This creates spin-polarized triplets that can engender a magnetization in the binormal direction, and plays the role typically provided by a magnetic field. We therefore see that strain can be an effective generator and controller/switcher of spin-current and magnetization in SNS junctions.
IV Discussion and Conclusion
We showed in Section III.1 that bending-strain in an isolated superconducting film induces p-wave triplet pairings throughout the superconductor for uniform curvature. Their magnitude is always around one order lower than the s-wave singlet, and die out as the parent correlation diminishes. Since the superconducting order parameter started out fully s-wave singlet, our self-consistency calculations included only the s-wave gap. We expect that if the p-wave correlations are included in the self-consistency requirement, p-wave pairings may stabilize to form the majority pairing population for increasingly high strain. For non-constant curvatures in the film, the gap would reduce locally in areas with high strain, and the p-wave contribution would increase only locally. We would also expect the sign of the curvature in each arm of the SNS junction to affect the ground state Skarpeid et al. (2024). For an $S$-shaped junction with opposite but equal curvatures in each arm, we would expect the spin magnetization to become vanishingly small as the spin polarization contributions from each interface would cancel Skarpeid et al. (2024).
Bending-strain translates into an effective SOC with symmetry breaking axis in the radial direction. Since the arc length defining the radius of curvature changes for every layer in the film, the total bending-strain of any one sample depends on the film thickness according to Eq. 9. We showed results for thicknesses $N_{n}=(3,6,9)$, and for equal strains these thicknesses have both a different curvature and a different corresponding effective SOC. We provide an overview of their comparative values in Table C.1, were we for convenience also compare the SOC energy with respect to the superconducting gap. When combining the radial bending-strain with interfacial-Rashba SOC in thin-film heterostructures, these two different symmetry-breaking axes combine to give a nuanced mixture of odd- and even-frequency correlations. They can be harnessed in conjunction with the spontaneous charge current that exists in an SNS Josephson junction, giving pure spin-current even at zero phase difference in the junction. The spin current can drive a switchable magnetization in the normal metal. Strain alone may therefore also induce a superconducting diode effect Mao et al. (2024).
In further development of this setup, it would be possible to include intrinsic Dresselhaus SOC in the normal metal, as well as out-of-plane symmetry breaking Rashba SOC induced via a substrate. Furthermore, since the normal metal undergoes no in-situ deformation, it would also be possible to include extrinsic Rashba SOC from gating, giving an additional control over the interconversion mechanisms for the superconducting correlations. While we also restricted the analysis in this study to conventional superconductors, a natural extension would be to consider unconventional materials, or multiband superconductors. In those cases modifications to the electron-phonon interaction may play an even bigger role, and strain-deformation of the bands in two-band superconductors could have profound implications for stabilizing superconductivity at high magnetic fields Salamone et al. (2023).
Acknowledgements.
@@
Appendix A Pair amplitudes
The pairing amplitudes can be found by inserting the operators in Eq. (21) into (22)-(26). This gives the following expressions for the even- and odd-frequency s- and p-wave singlet and triplet amplitudes, and their derivatives
$\mathcal{S}_{0,i}$ $=$ $\frac{1}{2N_{b}}\sum_{n,k}^{{}^{\prime}}\Big{[}x^{*}_{i,n,k}u_{i,n,k}-w^{*}_{i,n,k}v_{i,n,k}\Big{]}\tanh(\beta\mathcal{E}_{n,k}),$ (33)
$\mathcal{S}_{i,\uparrow\downarrow}(\tau)$ $=$ $\frac{1}{2N_{b}}\sum_{n,k}^{{}^{\prime}}\bigl{(}x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\bigr{)}\bigl{(}f(2\mathcal{E}_{n,k})e^{2i\mathcal{E}_{n,k}\tau}+f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}\bigr{)},$ (34)
$\dot{\mathcal{S}}_{i,\uparrow\downarrow}(\tau=0)$ $=$ $\frac{-i}{N_{b}}\sum_{n,k}^{{}^{\prime}}\mathcal{E}_{n,k}\bigl{(}x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\bigr{)}\tanh(\beta\mathcal{E}_{n,k}),$ (35)
$\mathcal{P}_{i,0}^{s}(\tau)$ $=$ $\frac{1}{2N_{b}}\sum_{\pm,n,k}^{{}^{\prime}}\pm\bigg{[}\big{(}x^{*}_{i\pm 1,n,k}u_{i,n,k}-w^{*}_{i\pm 1,n,k}v_{i,n,k}\big{)}f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}$ (36)

### v2

@@ IV.3 Discretization and diagonalization
$-\frac{i}{2}\sum_{\langle\bm{i},\bm{j}\rangle,\sigma,\sigma^{\prime}}\alpha_{N}\>\hat{\mathcal{N}}(s)\cdot\left(\bm{\sigma}(s)\times\bm{d_{ij}}\right){c^{\dagger}_{\bm{i},\sigma}c_{\bm{j},\sigma^{\prime}},}$
@@
IV.4 Numerical details
In Fig. 6, we show the dimensions used in the numerical scheme: the total number of lattice sites in the tangential, normal, and binormal direction are respectively denoted $N_{s},N_{n}$, and $N_{b}$, while the number of lattice sites in the superconductors and normal metal individually in the tangential and normal direction are respectively denoted $N_{S}$ and $N_{N}$.
@@ IV.4 Numerical details
We initialize a superconductor(s) with an s-wave gap on a square lattice, which is solved self-consistently with an iterative diagonalization by updating the gap in the Hamiltonian [Eq. 27] with the expression in Eq. 6 until the absolute difference between iterations falls below a threshold, here set to $10^{-6}$. We set the number of sites in the binormal and tangential direction to be square, with $N_{b}=N_{s}=99$, and Fourier transform the binormal direction with periodic boundary conditions. The numerical implementation is simplified by selecting an odd number of lattice sites, since the end of the Brillouin zone does not contribute in this case because the binomial momentum is restricted by $ka=2\pi m/N_{b}$, where $m$ is an integer [in effect, we can ignore the last term in Eq. 32]. While a specific number is chosen for illustration, the results are not sensitive to the number of lattice sites and remain consistent across different configurations, although a more sophisticated model would be expected to show atomic-scale oscillation of the strength of the superconductivity Guo et al. (2004); Ummarino and Zaccone (2024).
For the results presented in Section II, we normalize energy scales to the hopping parameter $t$, and set the attractive potential parameter to $U=1.5$, the chemical potential to $\mu=1$, and the temperature to $T=0.2$. Furthermore, we normalize the curvature to the superconducting coherence length $\alpha_{N}=\kappa\xi$, with proportionality constant $a_{N}=1$. This can be used to relate the effective spin-orbit strength to the strain via $\alpha_{N}=\epsilon_{ss}\xi/n$, where $n=(N_{n}-1)a/2$ is the length from the center of the film to the outermost layer (see Fig. 1), and $a$ is the lattice spacing. Using this, any specific strain values indicated in Section II refer to the strain at the upper/lower plane of the superconductor. For illustrative purposes, we selected parameters corresponding to superconducting niobium, with $\alpha_{N}=2.3\epsilon_{ss}/(N_{n}-1)$, where $\epsilon_{ss}$ is the strain in percent. The choice of material is arbitrary, and was chosen only to provide a familiar scale in comparison with purely spin-orbit-coupled systems, and the results that were presented will apply to any conventional thin-film superconductorBlatt and Thompson (1963); Guo et al. (2004); Chiang (2004); Ummarino and Zaccone (2024).
In the limiting case $N_{n}\to\infty$, the Rashba coefficient $\alpha_{N}\to 0$, suppressing p-wave pairings and spin-currents. Conversely, a strain-induced electric field $\mathbf{E}\propto\kappa\hat{\mathbf{\mathcal{N}}}$ requires $N_{n}\geq 2$. In order to emphasize the thickness dependence of pairing and spin responses under strain we considered the range $N_{n}\in(3,9)$, and set $\xi/a=2.3$ so that $\alpha_{N}$ falls within the common range for SOC-systems. The thin-film Hamiltonian is solved by summing over one-dimensional chains with a binormal momentum index $k$ and averaging over the first Brillouin zone, as described in Eqs. 26 and 32. Strain and curvature are incorporated as spin-orbit coupling, with the Pauli matrices rotating according to Eq. 16. The coordinate system is oriented such that the normal metal, when present, lies in the $y-z$ plane, as illustrated in Fig. 1.
Acknowledgments
@@
Supplementary Information:
Bending-strain effects in conventional superconductors and superconducting junctions
Kjell S. Heinrich,1 Henning G. Hugdal,1 Morten Amundsen,1 and Sol H. Jacobsen1
Center for Quantum Spintronics, Department of Physics, NTNU,
Norwegian University of Science and Technology, NO-7491 Trondheim, Norway
V Pair amplitudes
In the main text, adhering to the SPOT symmetry restrictions, we defined the s- and p-wave singlet and triplet amplitudes as
$\mathcal{S}_{\bm{i},0}$ $=\frac{1}{2}\big{[}\big{<}{c_{\bm{i},\uparrow}c_{\bm{i},\downarrow}}\big{>}-\big{<}{c_{\bm{i},\downarrow}c_{\bm{i},\uparrow}}\big{>}\big{]},$ (S1)
$\mathcal{S}_{\bm{i},\uparrow\downarrow}(\tau)$ $=\frac{1}{2}\big{[}\big{<}{c_{\bm{i},\uparrow}(\tau)\>c_{\bm{i},\downarrow}(0)}\big{>}+\big{<}{c_{\bm{i},\downarrow}(\tau)\>c_{\bm{i},\uparrow}(0)}\big{>}\big{]},$ (S2)
$\begin{split}\mathcal{P}_{\bm{i},0}^{\bm{m}}(\tau)&=\sum_{\pm}\!\pm\frac{1}{2}\big{[}\big{<}{c_{\bm{i},\uparrow}(\tau)\>c_{\bm{i}\pm\bm{m},\downarrow}(0)}\big{>}\\[-5.0pt] &\qquad\qquad-\big{<}{c_{\bm{i},\downarrow}(\tau)\>c_{\bm{i}\pm\bm{m},\uparrow}(0)}\big{>}\big{]},\end{split}$ (S3)
$\mathcal{P}^{\bm{m}}_{\bm{i},\uparrow\downarrow}$ =±±$=\sum_{\pm}\!\pm\frac{1}{2}\big{[}\big{<}{c_{\bm{i},\uparrow}c_{\bm{i}\pm\bm{m},\downarrow}}\big{>}+\big{<}{c_{\bm{i},\downarrow}c_{\bm{i}\pm\bm{m},\uparrow}}\big{>}\big{]},$ (S4)
$\mathcal{P}^{\bm{m}}_{\bm{i},\sigma\sigma}$ =±±$=\sum_{\pm}\!\pm\big{<}{c_{\bm{i},\sigma}c_{\bm{i}\pm\bm{m},\sigma}}\big{>},$ (S5)
where $\bm{m}$ can be $a\hat{\mathcal{T}}(s)$ or $a\hat{\mathcal{B}}(s)$ – one lattice spacing in the tangential or binormal direction, respectively. The annihilation and creation operators can be expressed in terms of the Bogoliubov quasiparticle operators
$c_{i,k,\uparrow}$ $=\sum_{n}u_{i,n,k}\gamma_{n,k},$ (S6)
$c_{i,k,\downarrow}$ $=\sum_{n}v_{i,n,k}\gamma_{n,k},$ (S7)
$c^{\dagger}_{i,-k,\uparrow}$ $=\sum_{n}w_{i,n,k}\gamma_{n,k},$ (S8)
$c^{\dagger}_{i,-k,\downarrow}$ $=\sum_{n}x_{i,n,k}\gamma_{n,k}.$ (S9)
The pairing amplitudes can be found by inserting the operators in Eq. (S6)-(S9) into (S1)-(S5). This gives the following expressions for the even- and odd-frequency s- and p-wave singlet and triplet amplitudes, and their derivatives
$\mathcal{S}_{0,i}$ $=$ $\frac{1}{2N_{b}}\sum_{n,k}^{{}^{\prime}}\Big{[}x^{*}_{i,n,k}u_{i,n,k}-w^{*}_{i,n,k}v_{i,n,k}\Big{]}\tanh(\beta\mathcal{E}_{n,k}),$ (S10)
$\mathcal{S}_{i,\uparrow\downarrow}(\tau)$ $=$ $\frac{1}{2N_{b}}\sum_{n,k}^{{}^{\prime}}\bigl{(}x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\bigr{)}\bigl{(}f(2\mathcal{E}_{n,k})e^{2i\mathcal{E}_{n,k}\tau}+f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}\bigr{)},$ (S11)
$\dot{\mathcal{S}}_{i,\uparrow\downarrow}(\tau=0)$ $=$ $\frac{-i}{N_{b}}\sum_{n,k}^{{}^{\prime}}\mathcal{E}_{n,k}\bigl{(}x^{*}_{i,n,k}u_{i,n,k}+w^{*}_{i,n,k}v_{i,n,k}\bigr{)}\tanh(\beta\mathcal{E}_{n,k}),$ (S12)
$\mathcal{P}_{i,0}^{s}(\tau)$ $=$ $\frac{1}{2N_{b}}\sum_{\pm,n,k}^{{}^{\prime}}\pm\bigg{[}\big{(}x^{*}_{i\pm 1,n,k}u_{i,n,k}-w^{*}_{i\pm 1,n,k}v_{i,n,k}\big{)}f(-2\mathcal{E}_{n,k})e^{-2i\mathcal{E}_{n,k}\tau}$ (S13)